OpenStudy (anonymous):

Need help with Trigonometric Identies cos^2x - sin^2x = 2cos^2x - 1 If someone could also explain the process it would be a big help

OpenStudy (anonymous):

Oh this one easy!! I am usually terribel at these

OpenStudy (anonymous):

Let us start with the Left hand side

OpenStudy (anonymous):

As we know Sin^2x+cos^2x=1

OpenStudy (anonymous):

So let us say in our original equation Cos^2x-sin^2x we subsitute the -sin^2x with something else so it easier to solve

OpenStudy (anonymous):

As we know from the identity that i showed you above -sin^2x=Cos^2x-1 so we will plug this into -sin^2x in our original equation

OpenStudy (anonymous):

cos^2x+(cos^2x-1) 2cos^2x-1

OpenStudy (anonymous):

And we got the same answer as the right hand side

OpenStudy (anonymous):

Did you get how i used the identity to sove it??? Cuz that was really not clear

OpenStudy (anonymous):

Haha not exactly, it was a bit hard to follow

OpenStudy (anonymous):

ok well i showed you that there is one identity Sin^2x+cos^2x=1 Do u know this?

OpenStudy (anonymous):

Are u there?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So u got that?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

ok now if I wld like i can rearrange this equation right?

OpenStudy (anonymous):

cos^2x=-sin^2x+1

OpenStudy (anonymous):

whtvr i can rearrange it which ever way i want as long as it equals the same

OpenStudy (anonymous):

Yup Cos^2x = 1 - Sin^2x or Sin^2x = 1 - Cos^2x

OpenStudy (anonymous):

Yahoo!!!!

OpenStudy (anonymous):

ok so I decide to reaarange it differently I bring the sin^2x to the right side and bring the one to the left side

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

cos^2x-1=-sin^2x

OpenStudy (anonymous):

I just flipped some values over that is it

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

So then since -sin^2x equals cos^2x1 I can plug this into the original equation

OpenStudy (anonymous):

cos^2x+(cos^2x-1)

OpenStudy (anonymous):

I follow what you're doing, but I'm confused on how to write that

OpenStudy (anonymous):

Well I just showed you how

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

Then open up the brackets and simplify and there is ur answer

OpenStudy (anonymous):

Oh I see, thank you! Would you mind helping me on one more? tanxcscxcosx = 1

OpenStudy (anonymous):

let me see if i can figure this out. I am terrible at this!!!!

OpenStudy (anonymous):

Oh this one i ssimple

OpenStudy (anonymous):

Oh really? Cause I was trying to do the problem and it came out retarded

OpenStudy (anonymous):